4. Twenty five high school students complete a preparation program for taking the SAT test. Here are the SAT scores from the 25 students who completed the SAT prep program: 434 694 457 534 720 400 484 478 610 641 425 636 454 514 563 370 499 640 501 625 612 471 598 509 531 The mean of these scores is 536.00. We know that the population average for SAT scores is 500 with a standard deviation of 100. The question is, are these students SAT scores (M = 536) significantly greater than a population mean of 500 with a population standard deviation of 100? Note that the creators of the SAT prep program claims that it will increase (and not decrease) your SAT score.
Answer #1

H0: Null hypothesis:mu leq 500

HA:Alternative hypothesis: μ> 500

SE = sigma /sqrt{n}

= 100/V25 = 4

Test statistic is:
Z = (536 - 500)/4 = 9

Take alpha =0.05

One Tail - Right Side Test

From Table, critical value of Z = 1.645

Since the calculated value of Z = 9 is greater than critical value of Z = 1.645,the difference is significant. Reject null hypothesis.

Conclusion:
The data support the claim that SAT prep program will increase SAT score.